OpenStudy (anonymous):

solve the system by the substitution method x-6y=-2 7x-5y=-14 a. no solution b. (-2,0) c. (-1,-2) d.( 2,-1)

myininaya (myininaya):

solve the first for x giving you: x=-2+6y now substitute this x in the second equation giving you: 7(-2+6y)-5y=-14 (This one can be solved for y)

myininaya (myininaya):

so the second equation is -14+42y-5y=-14 -14+37y=-14 now we add 14 on both sides giving us 37y=0 so y=0. Now we plug y=0 into our first equation giving us x=-2+6(0)=-2 so our solution is (-2,0)